(1 point) The region W is the cone shown below. The angle at the vertex is 1/2, and the top is flat and at a height of 3. WriCan someone help me with this?
Answer #1

First note that is *3 Conversion from cortesian to cylindricalis (x, y, z) to (8,0, 3) Putrs Sorry tan o = 4/x => dandyda ridfor spherical coordinates let x= psin & caso y = p sino sino = andy dz = gr sin odode dip za o cos where, s = dist from origiX z (a Cartesian is * an the ny plane is In the y z plane y=Z at kr/4 This angle st is 7/4 by symmetry = -2 Similarly, di xz2 =27 4* Я osz 23 So; 2² 8 - z & O= tail y It is clear from the projection on ny plane that Osrs z & o 20 sat a = bs 3 e- & V

The answer is in the pic. If any doubt still remained, let me know in the comment section.

If this solution helped, please don't forget to upvote to encourage us. We need your support. Thanks ☺☺☺ :)